Show that the vector v = (2, −5, 3) ∈ R 3

cannot be expressed as a linear combination of the

vectors v1= (1, −3, 2), v2= (2, −4, −1), v3= (1, −5, 7).


1
Expert's answer
2021-05-26T03:39:26-0400

Given,

The vector "\vec v=<2,-5,3>\ \ \in \R^3"

"\vec v_1=<1,-3,2>\\\vec v_2=<2,-4,-1>\\\vec v_3=<1,-5,7>"


Suppose "\vec v" can be expressed as linear combination of "\vec v_1,\vec v_2\ and \ \vec v_3"

Then,

"\vec v=x_1\vec v_1+x_2\vec v_2+x_2\vec v_3\\\vec v=x_1\begin{pmatrix}
 1 \\
 -3 \\
2
\end{pmatrix}+x_2\begin{pmatrix}
 2\\
 -4\\-1
\end{pmatrix}+x_3\begin{pmatrix}
 1\\
 -5\\7
\end{pmatrix}=\begin{pmatrix}
 2 \\
 -5\\3
\end{pmatrix}"


Augmented matrix of the above system is

"\begin{bmatrix}
 1&2&\ \ 1\ |\ \ \ \ \ 2 \\
 -3 & -4& -5\ |\ -5\\2&-1&7\ |\ \ \ 3
\end{bmatrix}"



Apply elementary row operation

"R_2\rightarrow R_2+3R_1\\R_3\rightarrow R_3-2R_1"


"=\begin{bmatrix}
 1&2&1&|2\\
 0&2&-2&|1\\0&-5&5&\ \ \ \ |-1
\end{bmatrix}"


Now,

"R_2\rightarrow \frac{1}{2}R_2\\\ \\R_3\rightarrow \frac{1}{5}R_3"


"=\begin{bmatrix}
 1&2&1&|2 \\
 0&1&-1&\ \ \ \ |1/2\\0&-1&1&\ \ \ \ \ \ \ \ |-1/5
\end{bmatrix}"


Also, "R_3\rightarrow R_3+R_2"

"=\begin{bmatrix}
 1&2&1&|2 \\
 0&1&-1&\ \ \ \ |1/2\\0&0&0&|3/10
\end{bmatrix}"


From above, we can conclude that there is pivot position in the augmented column.

Hence given system has no solution.

So, Vector "\vec v" can't be expressed as linear combination of "\vec v_1,\vec v_2\ and \ \vec v_3"